OpenStudy (anonymous):

Simplify the expression.(-1^5/-2^-3)^2

OpenStudy (anonymous):

Firstly (-1^5) will be -1 [-1 to the power of anything will always be 1 or -1] then -2^-3 this basically means 1/(-2^3) so now you have (-1 / (1/-2^3) which is the same as writing -1 * -2^3 work this out, and then square that answer Hope that helped

OpenStudy (anonymous):

wait, it has been a long time for me...I am lost on the squaring part

OpenStudy (anonymous):

my text book shows nothing to help, it confuses me

OpenStudy (anonymous):

Okay so to square something, you need to times it by itself So for example 4 squared = 4*4=16 5 squared = 5*5 = 25 6 squared = 6*6 = 36 So to square a number 'x' you need to do 'x*x' does that clarify?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

-1 * -2^3 = 8 then square that I get 64, am I right?

OpenStudy (anonymous):

yes that is correct :)

OpenStudy (anonymous):

thanks, I actually understand I a little better, I really appreciate it. I am going to save these steps for the other work...U ROCK!!!!

OpenStudy (anonymous):

No problem; happy to help! :)